[Hallc_running] [New Logentry] RC Update

hszumila at jlab.org hszumila at jlab.org
Mon Nov 8 09:35:02 EST 2021

Logentry Text:
3-pass energy was setup for Hall C yesterday during the day (https://logbooks.jlab.org/entry/3938406). We starting the swing shift with ion chamber calibrations. We note that the carbon hole/raster imaging looks a bit different than previously, but it wasn't clear if this is real or what the source of this is https://logbooks.jlab.org/entry/3938456.

The sieve runs were completed as 1-hour long runs as opposed to the 100k due to slow rate. The coincidence H(e,e'p) data was taken over night. There was a problem with the analyzing script so the statistics were not well-known until after we had moved on to the next setting. Both H(e,e'p) runs were taking with about 65% the desired statistics. 

We moved to production kinematics for low epsilon this morning, but beam is being tuned up to D right now, so we are standing by. 

There was an issue with one of the SHMS preshower voltages that Dave moved to another HV output connection and works fine now. It is not yet in the alarm handler. See https://logbooks.jlab.org/entry/3938576

We meet this morning at 10am via bluejeans: https://bluejeans.com/861439920

Agenda is here: https://hallcweb.jlab.org/wiki/index.php/RC_Meeting,_Monday,_November_8,_2021


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3938811
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20211108/17d548d5/attachment.html>

More information about the Hallc_running mailing list